OpenStudy (anonymous):

What is the domain of the function given in INTERVAL NOTATION:

OpenStudy (anonymous):

y=-3sec(pi-2x)+5

OpenStudy (anonymous):

we need to find the period and where the 1/costhteta =0

OpenStudy (anonymous):

its 2

OpenStudy (anonymous):

the period is 2

OpenStudy (anonymous):

i would disagree and say the period is pi shifted pi to the right and reflected

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but how do we go from there to interval notation

OpenStudy (anonymous):

\[\:[0,\:\frac{\pi }{4})\cup \left(\frac{\pi }{4},\:\frac{3\pi }{4}\right)\cup \left(\frac{3\pi }{4},\:\frac{5\pi }{4}\right)\cup \left(\frac{5\pi }{4},\:\frac{7\pi }{4}\right)\cup (\frac{7\pi }{4},\:2\pi ]\]

OpenStudy (anonymous):

is that the domain

OpenStudy (anonymous):

@mathstudent55 am I right now?

OpenStudy (anonymous):

@Deeezzzz am I right now?

OpenStudy (anonymous):

yes that is correct